OpenStudy (aristasiebrits8):

Hi! I REALLY need help in this one. It is super important. Part A: Solve C = (F - 32) for F. Part B: Determine the value of F when C = 8 °C. Part C: Solve -np - 80 > 60 for n. Show your work. Medal and Fan for the correct answer. Please, I REALLY need this.

OpenStudy (judygreeneyes):

Part A: Solve C = (F - 32) for F. Add 32 to both sides. C + 32 = F or F = C + 32 Part B: Determine the value of F when C = 8 °C. Using F = C + 32 plug in C = 8 F = 8 + 32 F = 40 Part C: Solve -np - 80 > 60 for n. Show your work. Not sure I understand this one completely, it doesn't look like it is related to the others. -np - 80 > 60 add 80 to both sides -np > 140 Divide both sides by p -n > 140/p Multiply both sides by -1 (need to change direction of > sign when you do this) n < -140/p

OpenStudy (usukidoll):

usually giving direct answers is against code of conduct but I'm not in the mood to have that enforced, so nice job :D

OpenStudy (judygreeneyes):

thx!

OpenStudy (aristasiebrits8):

Thanks so much for your help! You are an absolute life-saver!

OpenStudy (judygreeneyes):

You're welcome. I hope you understand how I solved it and can use that approach next time.

OpenStudy (skullpatrol):

Are you sure about the formula in part A? I think you missed a 5/9 $$\Huge C=\dfrac{5}{9}(F-32)$$

OpenStudy (judygreeneyes):

I agree with skullpatrol. If you are trying to convert from degrees C to degrees F or visa versa you need the scaling factor of 5/9.